Find the value of x if the arithmetic mean of 3 and 3x+5 is 8
Please help and explain how to
Thank you in advance

Answer by Alan3354(69443)

Find the value of x if the arithmetic mean of 3 and 3x+5 is 8
--------
The mean is the average.
(3 + 3x+5)/2 = 8
3x + 8 = 16
x = 8/3